# GLDPC-PC Codes for MIMO Systems with Iterative Detection and Decoding

### v1

@@ I Introduction
GLDPC [2] code generalizes low-density parity-check (LDPC) code [1] by replacing the single parity-check (SPC) constraints with more general constraints based on stronger component codes [3]. The belief propagation (BP) decoding of generalized low-density parity-check (GLDPC) codes iteratively pass extrinsic information between the variable nodes and check nodes in the Tanner graph where every CN represents a component code. The soft-input soft-output (SISO) decoder for the component code is necessary for BP decoding. One of the major obstacles in the development of GLDPC codes is the lack of low-complexity SISO decoders.
@@ III-C BP Decoding for GLDPC-PC
The use of the approximate equality sign here is due to insufficient iterations and cycles in $\mathbf{\Gamma}$. The decoder obtains $\hat{c}^{N}$ by hard decision based on $L_{i}^{\mathrm{APP}}$ and checks whether $\hat{c}^{N}$ is a codeword via whether the parity check of every component code is satisfied. The decoding terminates if $\hat{c}^{N}$ is a codeword or the number of iterations reaches $R$ and the $L_{i}^{\mathrm{APP}}$ at this time is the final soft output.
@@
The BP with SO-SCL decoder realizes the SISO decoder for GLDPC-PC codes. The computational complexity of BP with SO-SCL is at most $\mathcal{O}\left(RmNL\log_{2}N\right)$, where $N$ is the largest block length of all the component codes. Whereas the complexity of SO-SCL decoder for polar-like code is higher than the SISO decoder for SPC, the constraint of polar-like codes is more powerful than SPC. Therefore, less iterations are required for GLDPC-PC code to obtain similar performance with LDPC code and the complexity increase can be compensated by this.
@@ IV-C SISO MMSE-PIC Detection
$L_{i,j}^{\mathrm{E2}}=\frac{1}{\nu_{i}}\left(\min_{a\in\mathcal{Z}_{j}^{0}}\left|\tilde{x}_{i}-a\right|^{2}-\min_{a\in\mathcal{Z}_{j}^{1}}\left|\tilde{x}_{i}-a\right|^{2}\right).$ (23)
@@ V-B MIMO channel
Figure 4: BLER performance for (1024, 643) GLDPC-PC code and 5G LDPC code in $4\times 4$ MIMO channel. $N_{I}=R/D$ rounds of BP is performed once detection.
Figure 5: BLER performance for (1024, 643) GLDPC-PC code and 5G LDPC code in $8\times 8$ MIMO channel. $N_{I}=R/D$ rounds of BP is performed once detection.
@@ VI Conclusion
In this work, we first introduce the GLDPC-PC codes. These codes exhibit significantly lower error floors, which bring performance gains in the low BLER region in AWGN channel. The BP decoder with SO-SCL of GLDPC-PC is introduced, which adopts a low-complexity SO-SCL decoder for polar-like codes. We also adapt this BP with SO-SCL decoder in MIMO IDD systems coded by GLDPC-PC codes. Simulation results show the performance advantages of GLDPC-PC code with fewer rounds of decoding iteration. The GLDPC-PC code has the potential to perform better and exceed LDPC code with more elaborate construction.

### v2

@@ I Introduction
GLDPC [2] code generalizes low-density parity-check (LDPC) code [1] by replacing the single parity-check (SPC) constraints with more general constraints based on stronger component codes. The belief propagation (BP) decoding of generalized low-density parity-check (GLDPC) codes iteratively pass extrinsic information between the variable nodes and check nodes in the Tanner graph where every CN represents a component code. The soft-input soft-output (SISO) decoder for the component code is necessary for BP decoding. One of the major obstacles in the development of GLDPC codes is the lack of low-complexity SISO decoders.
@@ III-C BP Decoding for GLDPC-PC
The APP LLRs are only approximate values due to the cycles in $\mathbf{\Gamma}$. The decoder obtains $\hat{c}^{N}$ by hard decision based on $L_{i}^{\mathrm{APP}}$ and checks whether $\hat{c}^{N}$ is a codeword via whether the parity check of every component code is satisfied. The decoding terminates if $\hat{c}^{N}$ is a codeword or the number of iterations reaches $R$ and the $L_{i}^{\mathrm{APP}}$ at this time is the final soft output.
@@
The BP with SO-SCL decoder realizes the SISO decoding for GLDPC-PC codes. The computational complexity of BP with SO-SCL is at most $\mathcal{O}\left(RmNL\log_{2}N\right)$, where $N$ is the largest block length of all the component codes. Whereas the complexity of SO-SCL decoder for polar-like code is higher than the SISO decoder for SPC, the constraint of polar-like codes is more powerful than SPC. Therefore, less iterations are required for GLDPC-PC code to obtain similar performance with LDPC code and the complexity increase can be compensated by this.
@@ IV-C SISO MMSE-PIC Detection
$L_{i,j}^{\mathrm{E2}}=\frac{1}{\nu_{i}}\left(-\min_{a\in\mathcal{Z}_{j}^{0}}\left|\tilde{x}_{i}-a\right|^{2}+\min_{a\in\mathcal{Z}_{j}^{1}}\left|\tilde{x}_{i}-a\right|^{2}\right).$ (23)
@@ V-B MIMO Channel
Figure 4: BLER performance for (1024, 643) GLDPC-PC code and 5G LDPC code in $4\times 4$ MIMO channel. $N_{I}=R/D$ rounds of BP is performed after once detection.
Figure 5: BLER performance for (1024, 643) GLDPC-PC code and 5G LDPC code in $8\times 8$ MIMO channel. $N_{I}=R/D$ rounds of BP is performed after once detection.
@@ VI Conclusion
In this work, we first introduce the GLDPC-PC codes. These codes exhibit significantly lower error floors, which brings performance gains in the low BLER region in AWGN channel. The BP decoder is introduced for GLDPC-PC codes, which adopts the low-complexity SO-SCL decoder for polar-like codes. We also adapt this BP with SO-SCL decoder in MIMO IDD systems coded by GLDPC-PC codes. Simulation results show the performance advantages of GLDPC-PC code with fewer rounds of decoding iteration. The GLDPC-PC code has the potential to perform better and exceed LDPC code with more elaborate construction.